OpenStudy (anonymous):

Solve for x: sinx-cosx=√2

OpenStudy (cwrw238):

use the complementary angle method Rsinxcosa - Rcosxsina = Rsin(x-a) compare this with sinx - cos x = sqrt2 R cosa = 1 R sina = 1 so Rsina / Rcosa = tan a = 1 a = 45 degrees now you can find value of R and using Rsin(x - 45) = sqrt2 find x

OpenStudy (cwrw238):

oh - how to find R squaring R^2(cos^a + sin^a) = 2 R^2 = 2 and R = sqrt2

OpenStudy (anonymous):

sinx - cosx =sqrt(2) sin^2(x) -2cos(x)sin(x) + cos^2(x) = 2 sin^2(x)+cos^2(x) -2cos(x)sin(x)=2 1-2sinxcosx=2 -2sinxcosx=1 -2(0.5(sin(x+x)+sin(x-x))=1 using identities -sin2x+sin0=1 sin2x=-1 2x=-90 degrees + k360 x=-45 degrees + k180 or 2x=270 degrees +k360 x= 135 degrees + k180 where k is integer

OpenStudy (anonymous):

wow :d

OpenStudy (anonymous):

you could leave out -2(0.5(sin(x+x)+sin(x-x))=1 - that is an unnecessary step i did :) because sin2x=2sinxcosx is quite obvious already